Machine, weighing 800 kg; descends with the engine extinguished from a hill 40 m high. The driver presses the train pedal with his foot during the descent. At the top of the hill the car has a speed of 6 m / s while at the bottom of the hill 20 m / s. Find the amount of mechanical energy converted into heat while zoning the car.


1
Expert's answer
2021-05-23T16:34:59-0400
"Q=0.5m(v^2-V^2)+mgh\\Q=-0.5(800)(20^2-6^2)+(800)(9.8)(40)\\=168000\ J"